Solve the problem with complete solution. In an experiment to determine the coefficient of linear expansion for aluminum, a .50 m rod is heated from 20 degree Celsius to 100 degree Celsius. The length increases to .84 m. Find the coefficient of linear expansion for aluminum.


1
Expert's answer
2021-12-28T13:36:40-0500
"\Delta L=\alpha L_0\Delta T,""\alpha=\dfrac{L-L_0}{L_0\Delta T}=\dfrac{0.84\ m-0.5\ m}{0.5\ m\times80^{\circ}C}=8.5\times10^{-3}\ (^{\circ}C)^{-1}."